OpenStudy (kmullis6):

What is the measure of one angle of a regular octagon? If necessary, round to the nearest degree. Choices: a) 120 degrees. b) 135 degrees. c) 140 degrees. d) 150 degrees.

ganeshie8 (ganeshie8):

angle = (n-2)*180/n put n=8

OpenStudy (anonymous):

Or, exterior angle = 360/8 so interior is....

OpenStudy (kmullis6):

divide it?

OpenStudy (anonymous):

divide what?

OpenStudy (kmullis6):

360/8... sorry i don't really understand. /:

OpenStudy (anonymous):



OpenStudy (anonymous):

Yes, / means divide by

OpenStudy (kmullis6):

45?

OpenStudy (anonymous):

So now, what is the interior?

OpenStudy (anonymous):